# Improved Scaling Laws in Linear Regression via Data Reuse

### v1

@@ D.3 Fluctuation error under the source condition
We start the proof by establishing concentration bounds on some quantities that appear in the bounds in Lemma D.1 and D.4.
@@
$T_{\mathbf{B}}$ :=$:=\begin{cases}1+\log(1/\delta)/N+{{\sf t}(\delta)}\cdot({L_{\mathtt{eff}}}\gamma)^{2}\cdot(1+\log(1/\delta)/N)^{2}~{}~{}~{}\text{when}~{}~{}~{}M\leq N/2,\\ \mathbf{B}_{\mathbf{B}}\cdot(1+\log(1/\delta)/N)^{2}+1~{}~{}~{}\text{when}~{}~{}~{}M>N/2,\end{cases}$
@@
Therefore, substituting Eq. (22a)—(22c) and (24) into the expression in $a_{\max},\mathbf{B}_{\bm{\Delta}}$ and $\mathsf{F}$, applying Eq. (15) to bound $\widetilde{D}^{\mathsf{U}}$ (and $\operatorname{tr}({\bm{\widehat{\Sigma}}}^{1/\alpha})$ for some $\alpha=a+\varepsilon>a$), using part 2 of Lemma D.1444More specifically, we apply part 2 of Lemma D.1 on the event with probability at least $1-\exp(-\Omega(N))$ where conditions on ${\bm{\widehat{\Sigma}}}$ specified in Lemma E.8 hold, and apply part 1 of Lemma D.1 for some $\alpha=a+\varepsilon>1$ otherwise. and taking expectation w.r.t. $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ conditioned on $\mathbf{S}$ and $\mathbf{w}^{*}$, it can be verified that
@@
Setting $s=0$ in Eq. (22c), we have $B_{\bm{\xi}}\geq 1/2$ when $\gamma{L_{\mathtt{eff}}}B_{\bm{\xi}}/N\geq 1/3$, which happens with probability at least $1-N^{-c_{1}/c_{2}}$ for some constant $c_{1}>0$ when $\gamma\leq c_{2}/\log N$ for some $c_{2}>0$. Moreover, by the concentration properties on $\mu_{j}({\bm{\widehat{\Sigma}}})$ and $\mu_{j}({\bm{\Sigma}})$ in the previous discussion, the assumptions on $\gamma$, and a union bound, conditioned on $\mathbf{S}$ such that $\mu_{j}({\bm{\Sigma}})\eqsim j^{-a}$ for $j\in[M]$ (which happens with probability at least $1-e^{-\Omega(M)}$), we have with probability at least $1/2$ over the randomness of $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ that
$\mu_{j}({\bm{\widehat{\Sigma}}})$ $\eqsim j^{-a}~{}~{}~{}~{}\text{for}~{}~{}~{}~{}j\leq\min\{M,N/c\},\text{~{}~{}and}$
@@
Following the proof of Lemma B.1 (Eq. 10 and 11), it can be verified that, conditioned on $\mathbf{S}$ and $\mathbf{w}^{*}$,
[…]
@@
Adopt the shorthand $\mathsf{V}_{t}^{(-i)}$ for $\mathbf{I}-\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})$ and $R_{i}$ for $\|({\bm{\widehat{\Sigma}}}^{(-i)}+\lambda\mathbf{I})^{-1/2}({\bm{\Sigma}}^{(-i)}+\lambda\mathbf{I})^{1/2}\|^{2}$.
@@
$\widetilde{D}^{\mathsf{U}}_{i}$ :=#$:=\#\{j\in[M]:\widehat{\lambda}^{(-i)}_{j}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{j:\widehat{\lambda}^{(-i)}_{j}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}^{(-i)}_{j},$
and $\widehat{\lambda}^{(-i)}_{j}$ is the $j$-th largest eigenvalue of ${\bm{\widehat{\Sigma}}}^{(-i)}$. Here, step (ii) uses Eq. (16) and step (iii) uses the fact that $\widehat{\lambda}^{(-i)}_{j}\leq\widehat{\lambda}_{j}$ for all $j\in[M]$ since ${\bm{\widehat{\Sigma}}}^{(-i)}\preceq{\bm{\widehat{\Sigma}}}$.
@@
Denote the eigenvalues of ${\bm{\Sigma}}_{\bm{\nu}}$ by $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\ldots\geq\widehat{\lambda}_{N}$ (let $\widehat{\lambda}_{j}=0$ for $j>M$). Choose $\iota=({L_{\mathtt{eff}}}\gamma)^{1/a}$. When $M\leq N/\tilde{c}$, we have $\mu_{j}({\bm{\Sigma}}_{\bm{\nu}})\eqsim j^{-a}$ for $j\leq M$ and otherwise 00. When $M>N/\tilde{c}$, we have $\iota\leq N/\tilde{c}$, $\widehat{\lambda}_{j}\eqsim j^{-a}$ for $j\leq N/\tilde{c}$ and otherwise $\widehat{\lambda}_{j}\lesssim j^{-a}$ by monotonicity of $\widehat{\lambda}_{j}$. In both cases, we have $\widehat{\lambda}_{j}\eqsim j^{-a}$ (or $\widehat{\lambda}_{j}=0$) for $j\leq\iota$ and $\widehat{\lambda}_{j}\lesssim j^{-a}$ for $j>\iota$.
@@
It remains to bound $\max_{i\in[N],t\in[L]}\|\bm{\Delta}^{(-i)}_{t}\|^{2}_{{\bm{\Sigma}}^{s}}$. Adopt the shorthand notation $a^{(-i)}_{t}:=y_{i}+\mathbf{x}_{i}^{\top}\mathbf{S}^{\top}\bm{\theta}^{(-i)}_{t-1}$ and recell $a_{\max}=\max_{i\in[N],t\in[L]}|a^{(-i)}_{t}|$. By taking the difference between the (GD) process and the (LOO-GD) process, we have
@@
Adopt the shorthand notation $\mathsf{V}_{t}=\mathbf{I}-\prod_{i=1}^{t}(\mathbf{I}-\gamma_{i}{\bm{\widehat{\Sigma}}})$. Choosing $\lambda=1/({L_{\mathtt{eff}}}\gamma)$ in the last display and taking the supremum over $t\in[L],i\in[N]$, we obtain
@@ E.1 General concentration bounds
[…]
for $\text{for every }j\leq M,\quad\left|\tilde{\lambda}_{j}-\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)\right|\leq c\left(\sqrt{\frac{k}{M}}\lambda_{j}+\lambda_{k+1}+\sqrt{\frac{\sum_{i>k}\lambda_{i}^{2}}{M}}\right).$
@@
for $\text{for every }j\leq M,\quad\left|\tilde{\lambda}_{j}-\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)\right|\leq\frac{1}{2}\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)+c_{1}\lambda_{k+1},$
@@
for $\text{for every }j\leq k,\quad\left|\mu_{j}\left(\mathbf{S}_{0:k}\mathbf{H}_{0:k}\mathbf{S}_{0:k}^{\top}\right)-\lambda_{j}\right|\lesssim\sqrt{\frac{k+\log(1/\delta)}{M}}\lambda_{j}.$
@@
for $\text{for every }j\leq k,\quad\left|\mu_{j}\left(\mathbf{S}_{0:k}\mathbf{H}_{0:k}\mathbf{S}_{0:k}^{\top}\right)-\lambda_{j}\right|\lesssim\sqrt{\frac{k}{M}}\lambda_{j}.$
@@ E.2 Concentration bounds under power-law spectrum
Lemma E.8 (Eigenvalues of ${\bm{\widehat{\Sigma}}}$ under power-law spectrum).
@@
[…]

### v2

@@ D.3 Fluctuation error under the source condition
The proof follows from instantiating Lemma D.1 and D.4 under the source condtion. We start by establishing concentration bounds on some quantities that appear in the bounds in Lemma D.1 and D.4.
@@
$T_{\mathbf{B}}$ $:=\begin{cases}1+\log(1/\delta)/N+{{\sf t}(\delta)}\cdot({L_{\mathtt{eff}}}\gamma)^{2}\cdot(1+\log(1/\delta)/N)^{2}~~~\text{when}~~~M\leq N/2,\\ \mathbf{B}_{\mathbf{B}}\cdot(1+\log(1/\delta)/N)^{2}+1~~~\text{when}~~~M>N/2,\end{cases}$
@@
Therefore, substituting Eq. (24a)—(24c) and (26) into the expression in $a_{\max},\mathbf{B}_{\bm{\Delta}}$ and $\mathsf{F}$, applying Eq. (16) to bound $\widetilde{D}^{\mathsf{U}}$ (and $\operatorname{tr}({\bm{\widehat{\Sigma}}}^{1/\alpha})$ for some $\alpha=a+\varepsilon>a$), using part 2 of Lemma D.1444More specifically, we apply part 2 of Lemma D.1 on the event with probability at least $1-\exp(-\Omega(N))$ where conditions on $^{\bm{\widehat{\Sigma}}}$ specified in Lemma E.8 hold, and apply part 1 of Lemma D.1 for some $\alpha=a+\varepsilon>1$ otherwise. and taking expectation w.r.t. $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ conditioned on $\mathbf{S}$ and $\mathbf{w}^{*}$, it can be verified that
@@
[…]
@@
Following the proof of Lemma B.1 (Eq. 11 and 12), it can be verified that, conditioned on $\mathbf{S}$ and $\mathbf{w}^{*}$,
$\|\prod_{i=1}^{t}(\mathbf{I}-\gamma_{i}{\bm{\widehat{\Sigma}}}^{(-i)})\mathbf{v}^{*}\|^{2}_{{\bm{\Sigma}}}$ $\lesssim\begin{cases}B_{\sf F,1}\overset{d}{=}\|\mathbf{Z}\widebar{\mathbf{w}}^{*}\|_{2}^{2}+({L_{\mathtt{eff}}}\gamma)^{2}\|\mathbf{Z}^{\top}\mathbf{Z}\|_{2}^{2}\bm{1}_{\{\mathbf{Z}^{\top}\mathbf{Z}\nsucceq\mathbf{I}_{M}/5\}}\cdot\|\widebar{\mathbf{w}}^{*}\|_{2}^{2}~~~\text{when}~~~M\leq N/2,\\ B_{\sf F,2}\overset{d}{=}B_{\sf F,3}\cdot\|\mathbf{Z}\widebar{\mathbf{w}}^{*}\|_{2}^{2}+\|\widebar{\mathbf{w}}^{*}\|_{2}^{2}~~~\text{when}~~~M>N/2,\end{cases}$
@@
[…]
@@
$\widetilde{D}^{\mathsf{U}}_{i}$ $:=\#\{j\in[M]:\widehat{\lambda}^{(-i)}_{j}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{j:\widehat{\lambda}^{(-i)}_{j}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}^{(-i)}_{j}$,
and $\widehat{\lambda}^{(-i)}_{j}$ is the $j$-th largest eigenvalue of ${\bm{\widehat{\Sigma}}}^{(-i)}$. Here, step (ii) uses Eq. (17) and step (iii) uses the fact that $\widehat{\lambda}^{(-i)}_{j}\leq\widehat{\lambda}_{j}$ for all $j\in[M]$ since $^{\bm{\widehat{\Sigma}}}^{(-i)}\preceq{\bm{\widehat{\Sigma}}}$.
@@
Denote the eigenvalues of ${\bm{\Sigma}}_{\bm{\nu}}$ by $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\ldots\geq\widehat{\lambda}_{N}$ (let $\widehat{\lambda}_{j}=0$ for $j>M$). Choose $\iota=({L_{\mathtt{eff}}}\gamma)^{1/a}$. When $M\leq N/\tilde{c}$, we have $\mu_{j}({\bm{\Sigma}}_{\bm{\nu}})\eqsim j^{-a}$ for $j\leq M$ and otherwise 0. When $M>N/\tilde{c}$, we have $\iota\leq N/\tilde{c}$, $\widehat{\lambda}_{j}\eqsim j^{-a}$ for $j\leq N/\tilde{c}$ and otherwise $\widehat{\lambda}_{j}\lesssim j^{-a}$ by monotonicity of $\widehat{\lambda}_{j}$. In both cases, we have $\widehat{\lambda}_{j}\eqsim j^{-a}$ (or $\widehat{\lambda}_{j}=0$) for $j\leq\iota$ and $\widehat{\lambda}_{j}\lesssim j^{-a}$ for $j>\iota$.
@@
[…]
@@
[…]
@@ E.1 General concentration bounds
Let $\mathbf{S}\in\mathbb{R}^{M\times d}$ be a random sketching matrix with i.i.d. entries $\mathbf{S}_{ij}\sim\mathcal{N}(0,1/M)$.555dd can be $\infty$. Then there exists some absolute constant $c>1$ such that for any $M\geq 1$ and $0\leq k\leq M$, with probability at least $1-e^{-\Omega(M)}-e^{-\Omega(k)}$, we have
$\text{for every }j\leq M,\quad\left|\tilde{\lambda}_{j}-\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)\right|\leq c\left(\sqrt{\frac{k}{M}}\lambda_{j}+\lambda_{k+1}+\sqrt{\frac{\sum_{i>k}\lambda_{i}^{2}}{M}}\right).$
@@
$\text{for every }j\leq M,\quad\left|\tilde{\lambda}_{j}-\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)\right|\leq\frac{1}{2}\left(\lambda_{j}+\frac{\sum_{i>k}\lambda_{i}}{M}\right)+c_{1}\lambda_{k+1},$
@@
$\text{for every }j\leq k,\quad\left|\mu_{j}\left(\mathbf{S}_{0:k}\mathbf{H}_{0:k}\mathbf{S}_{0:k}^{\top}\right)-\lambda_{j}\right|\lesssim\sqrt{\frac{k+\log(1/\delta)}{M}}\lambda_{j}.$
@@
$\text{for every }j\leq k,\quad\left|\mu_{j}\left(\mathbf{S}_{0:k}\mathbf{H}_{0:k}\mathbf{S}_{0:k}^{\top}\right)-\lambda_{j}\right|\lesssim\sqrt{\frac{k}{M}}\lambda_{j}.$
@@ E.2 Concentration bounds under power-law spectrum
Lemma E.8 (Eigenvalues of $^{\bm{\widehat{\Sigma}}}$ under power-law spectrum).
@@
Note that $\mathbf{S}\mathbf{X}^{\top}/\sqrt{N}\overset{d}{=}{\bm{\Sigma}}^{1/2}\mathbf{Z}^{\top}$, where $\mathbf{Z}\in{\mathbb{R}}^{M\times N}$ has i.i.d. entries $\mathbf{Z}_{ij}\sim\mathcal{N}(0,1/N)$ conditioned on $\mathbf{S}$. Thus, $\mu_{j}({\bm{\widehat{\Sigma}}})=\mu_{j}(\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top})$ for $j\leq\min\{M,N\}$. Let $(\widehat{\lambda}_{i})_{i=1}^{N}$ denote the eigenvalues of $\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top}$ in non-increasing order. Using Lemma E.2 with $k=N/c$ for some sufficiently large constant $c$ and noting that $\sum_{i>k}i^{-{a}}\lesssim k^{1-{a}}$, we have